# LLM Watermark Evasion via Bias Inversion

### v3

@@ Abstract
Watermarking for large language models (LLMs) embeds a statistical signal during generation to enable detection of model-produced text. While watermarking has proven effective in benign settings, its robustness under adversarial evasion remains contested. To advance a rigorous understanding and evaluation of such vulnerabilities, we propose the Bias-Inversion Rewriting Attack (BIRA), which is theoretically motivated and model-agnostic. BIRA weakens the watermark signal by suppressing the logits of likely watermarked tokens during LLM-based rewriting, without any knowledge of the underlying watermarking scheme. Across recent watermarking methods, BIRA achieves over 99% evasion while preserving the semantic content of the original text. Beyond demonstrating an attack, our results reveal a systematic vulnerability, emphasizing the need for stress testing and robust defenses.
@@ 1 Introduction
The rapid advancement and proliferation of large language models (LLMs) (Minaee et al., 2024; Wang et al., 2024a) have intensified concerns about their misuse, ranging from the spread of misleading content (Monteith et al., 2024; Wang et al., 2024b; Papageorgiou et al., 2024) to threats to academic integrity, such as cheating (Stokel-Walker, 2022; Kamalov et al., 2023). To address these risks, watermarking has been proposed as a promising approach for detecting LLM-generated content (Aaronson and Kirchner, 2022; Kirchenbauer et al., 2024b). The core idea is to embed an imperceptible statistical signal into generated text, for example, by partitioning the vocabulary into “green” and “red” lists using a secret key, and biasing generation toward the green list. A detector then identifies LLM-generated text by checking for statistical overrepresentation of green tokens.
Recent work shows that watermarking is robust against common evasion strategies, such as text insertion, text substitution, and text deletion (Kirchenbauer et al., 2024b; Liu et al., 2024; Zhao et al., 2024; Lee et al., 2024; Lu et al., 2024). This robustness has drawn significant attention and spurred movement toward deployment. For instance, OpenAI has discussed adding watermarking to its products (Bartz and Hu, 2023), and U.S. policymakers have proposed legislation requiring watermarks for AI-generated content (Tong, 2024).
@@
However, recent studies (Raffel et al., 2020; Cheng et al., 2025; Wu and Chandrasekaran, 2024; Chen et al., 2024; Jovanović et al., 2024; Diaa et al., 2024) have questioned the robustness of watermarking, noting that existing methods have not been sufficiently stress-tested and showing that watermarks can be evaded through sophisticated strategies. These approaches fall into two categories: query-based attacks (Wu and Chandrasekaran, 2024; Chen et al., 2024; Jovanović et al., 2024), which focus on query strategies to recover the green token set but require unrestricted access to the target model via repeated queries and often rely on knowledge of the watermarking scheme, and query-free attacks (Raffel et al., 2020; Cheng et al., 2025), typically based on paraphrasing, which avoid this strong assumption but achieve only limited attack success and often distort semantic meaning.
In response to these limitations and to advance the understanding of watermarking vulnerabilities, we propose the Bias-Inversion Rewriting Attack (BIRA), motivated by a theoretical analysis of green–red list watermarking showing that reducing the probability of generating green tokens by $\delta>0$ during the rewriting of watermarked text causes the overall detection probability to decay exponentially in $\delta$. To achieve this, BIRA applies a negative bias to a proxy set of green tokens during paraphrasing with an LLM, without requiring knowledge of the underlying scheme. It consistently evades detection across a wide range of recent watermarking algorithms while preserving the semantics of the original text.
@@
  • We provide a theoretical characterization of a fundamental vulnerability of current green-red list watermarking schemes.
  • Building on this theoretical insight, we introduce BIRA, a query-free attack that weakens the watermark signal by applying a negative logit bias to likely watermarked tokens.
  • We conduct extensive experiments demonstrating that BIRA achieves state-of-the-art evasion rates against recent watermarking algorithms while maintaining high semantic fidelity.
@@ 2 Related Work
LLM watermarking. Kirchenbauer et al. (2024a) introduced a widely used green–red list watermarking scheme that partitions the vocabulary into green and red sets and embeds a detectable statistical signal by adding a positive logit bias to green tokens. Subsequent studies have enhanced its robustness by improving key generation and detection (Kirchenbauer et al., 2024b; Liu et al., 2023a; Zhao et al., 2024; Liu et al., 2024; Lee et al., 2024; Lu et al., 2024) or by preserving the original LLM distribution (Wu et al., 2023). Other lines of work investigate sampling-based watermarking approaches (Aaronson and Kirchner, 2022; Hu et al., 2023; Christ et al., 2024).
Watermark evasion attacks. Watermark evasion attacks can be broadly categorized under the threat model into two types: query-based and query-free. Query-based attacks (Jovanović et al., 2024; Chen et al., 2024; Wu and Chandrasekaran, 2024) identify the green token set by issuing a large number of carefully crafted prefix prompts. While effective, this strategy is computationally expensive and has a limited practicality in rate-limited API settings. Additionally, these methods often require knowledge of the watermarking scheme (Wu and Chandrasekaran, 2024) or access to top-$k$ token probabilities of the watermarked model (Chen et al., 2024), which is typically infeasible for proprietary systems.
In contrast, query-free attacks (Kirchenbauer et al., 2024a; Krishna et al., 2023; Cheng et al., 2025; Diaa et al., 2024) avoid these assumptions by operating directly on generated text without interacting with the watermarked model, typically relying on paraphrasing to obscure the statistical signal. Early work (Kirchenbauer et al., 2024a) introduced simple transformation attacks, such as inserting emojis or human-written fragments into watermarked text. More advanced methods either fine-tune an LLM as a paraphrasing expert (Krishna et al., 2023; Diaa et al., 2024) or use a masking-and-rewriting strategy that targets high-entropy tokens and regenerates them with an LLM (Cheng et al., 2025). However, these post-processing approaches often achieve limited attack success, can distort the original meaning, and in some cases require training an additional model. In contrast, our method attains a substantially higher attack success rate while preserving semantic fidelity, without training any extra model in the fully black box scenarios.
@@ 3 Preliminary
Figure 1: Illustration of BIRA. A watermarked LLM typically increases the likelihood of sampling green tokens by adding a positive bias $\gamma>0$ to their logits at each generation step. In contrast, BIRA applies a negative bias $\beta<0$ to a proxy set of green tokens (since the true set is unknown), thereby suppressing their sampling probability. This inversion lowers the probability of generating green tokens and weakens the watermark signal, enabling the paraphrased text to evade detection.
@@ 4 Method
In this section, we first present the theoretical analysis of watermarking vulnerabilities that our attack exploits (Section 4.1), and then describe the attack algorithm (Section 4.2). Figure 1 provides an overview of BIRA.
4.1 Theoretical Analysis of Watermarking Vulnerabilities
The goal of a watermark evasion attack is to diminish the overrepresentation of green tokens in a text to a level that cannot be detected statistically. We first show that common watermark detectors, which rely on test statistics like the z-score, are functionally equivalent to a simple threshold test on the empirical green token rate, $\hat{p}(y;\mathcal{W}_{k})$ (Theorem 1). Building on this, we prove that if the average probability of generating a green token across the sequence stays below the detection threshold by a margin $\delta$, then the detection probability decreases exponentially in $\delta$ (Theorem 2).
Theorem 1.
@@ 4.1 Theoretical Analysis of Watermarking Vulnerabilities
$Z(y;\mathcal{W}_{k})=h\!\left(\hat{p}(y;\mathcal{W}_{k})\right),\qquad\hat{p}(y;\mathcal{W}_{k})=\frac{1}{N}\sum_{n=0}^{N-1}\mathbf{1}\{y^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\},$
@@
Theorem 1 shows that detection can be expressed in terms of the empirical green rate $\hat{p}(y;\mathcal{W}_{k})$ with threshold $p_{\tau}$. Using this, we now demonstrate that suppressing the average green token probability across the sequence yields exponential decay in the detection probability (Theorem 2).
Theorem 2.
Let $\tilde{y}=[\tilde{y}^{(0)},\ldots,\tilde{y}^{(N-1)}]$ be the attacker’s output and let $p_{\tau}$ be the detection threshold. If there exists $\delta>0$ such that the average conditional green-token probability satisfies
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\}\,\middle|\,\tilde{y}^{0:n-1}\right]\;\leq\;p_{\tau}-\delta,$
@@
$\Pr\!\big[\mathcal{D}(\tilde{y},\mathcal{W}_{k})=1\big]\;\leq\;\exp\!\left(-\frac{1}{2}\,N\,\delta^{2}\right),$
Theorem 2 shows that if the average probability of sampling a green token stays at least $\delta$ below the detector threshold $p_{\tau}$, then the detection probability decays exponentially in $\delta$. In other words, even a small reduction in the green-token probability, when achieved on average over the sequence, is sufficient to make the text statistically undetectable and drive the overall detection probability toward zero. Proofs of Theorems 1 and 2 are provided in Appendix A. Although Theorem 2 is stated for token-level green–red list watermarking, the same analysis extends to recent sentence-level watermarking schemes. The details of the proof and additional experimental results are provided in Appendix E.
@@
Algorithm 1 Pseudocode for Bias-Inversion Rewriting Attack
1:System prompt $S$; Watermarked text $\hat{y}^{0:N-1}$; Language model $\mathcal{M}$; Percentile $q\in[0,1)$; Initial bias $\beta_{0}<0$; $\mathrm{lr}>0$; Max restarts $R$; Max length $L_{\text{max}}$; Window size $h$; threshold $\rho\in(0,1]$.
[…]
3:Compute self-information $I^{(n)}$ for each token $\hat{y}^{(n)}$ using the language model $\mathcal{M}$:
4:for $n=0,\ldots,N-1$ do
5:  $I^{(n)}\leftarrow-\log P_{\mathcal{M}}(\hat{y}^{(n)}|\hat{y}^{0:n-1})$
@@
21:   $\beta\leftarrow\min(0,\beta+\mathrm{lr})$ $\triangleright$ Reduce the strength of bias and restart
@@
24:   return $\tilde{y} \triangleright$ Return text
@@ 4.2 Bias-Inversion Rewriting Attack
As established in Theorem 2, successful watermark evasion requires suppressing green token generation across the sequence. In a black box setting, the adversary lacks access to the true green sets $\mathcal{G}(\mathcal{W}_{k})$, so we approximate them with a proxy set $\widehat{\mathcal{G}}$. Since watermarking schemes typically embed their signal in low-probability tokens, and token self-information (surprisal) is effective for detecting them Cheng et al. (2025), we identify such tokens using their self-information under a public language model $\mathcal{M}$:
$I^{(n)}=-\log P_{\mathcal{M}}(\hat{y}^{(n)}\mid\hat{y}^{0:n-1}).$
Given a watermarked text $\hat{y}=[\hat{y}^{(0)},\ldots,\hat{y}^{(N-1)}]$, let $\eta$ be the $q$th percentile of $\{I^{(n)}\}_{n=0}^{N-1}$. The proxy green set is
@@
where $\mathrm{id}(\cdot)$ maps a token to its vocabulary index. During paraphrase generation $\tilde{y}$, we add a negative logit bias $\beta<0$ to all tokens in $\widehat{\mathcal{G}}$ at each step $n$:
$l^{(n)}_{u}\leftarrow l^{(n)}_{u}+\beta\,\mathbf{1}\{u\in\widehat{\mathcal{G}}\}$.
This negative bias suppresses the generation of likely green tokens, which lowers the empirical green rate and thus enables evasion of detection. Since we rely on a proxy green set $\widehat{\mathcal{G}}$, some green tokens may be missed. However, Theorem 2 still holds as long as the average miss rate of the proxy set is bounded by $\epsilon$, i.e., $\frac{1}{N}\sum_{n=1}^{N}\Pr\left[\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\setminus\widehat{\mathcal{G}}\mid\tilde{y}^{0:n-1}\right]\leq\varepsilon$ and the average conditional probability of sampling from the proxy set is suppressed: $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\mathbf{1}\{\tilde{y}^{(n)}\in\widehat{\mathcal{G}}\}\mid\tilde{y}^{0:n-1}\right]\leq p^{\prime}_{\tau}-\delta$ with $p^{\prime}_{\tau}=p_{\tau}-\varepsilon$. Further details are provided in Appendix A.1.
Remark. Our approach differs from the masking-and-rewriting strategy of Cheng et al. (2025), which masks high-entropy tokens and then rewrites the masked spans. In contrast, we apply a negative logit bias to tokens in the proxy green set $\widehat{\mathcal{G}}$ at every decoding step. This consistently reduces the probability of sampling green tokens across the sequence and therefore lowers the detection probability, as established by Theorem 2. Moreover, by avoiding disruptive masking that breaks context, our method better preserves semantic fidelity.

### v4

@@ Abstract
Watermarking offers a promising solution for detecting LLM-generated content, yet its robustness under realistic query-free (black-box) evasion remains an open challenge. Existing query-free attacks often achieve limited success or severely distort semantic meaning. We bridge this gap by theoretically analyzing rewriting-based evasion, demonstrating that reducing the average conditional probability of sampling green tokens by a small margin causes the detection probability to decay exponentially. Guided by this insight, we propose the Bias-Inversion Rewriting Attack (BIRA), a practical query-free method that applies a negative logit bias to a proxy suppression set identified via token surprisal. Empirically, BIRA achieves state-of-the-art evasion rates ($99\%$) across diverse watermarking schemes while preserving semantic fidelity substantially better than prior baselines. Our findings reveal a fundamental vulnerability in current watermarking methods and highlight the need for rigorous stress tests.
@@ 1 Introduction
The rapid advancement and proliferation of large language models (LLMs) (Minaee et al., 2024; Wang et al., 2024a) have intensified concerns about misuse, ranging from the spread of misleading content (Monteith et al., 2024; Wang et al., 2024b; Papageorgiou et al., 2024) to threats to academic integrity such as cheating (Stokel-Walker, 2022; Kamalov et al., 2023). To address these risks, watermarking has been proposed as a promising approach for detecting LLM-generated content (Aaronson and Kirchner, 2022; Kirchenbauer et al., 2024b). The core idea is to embed an imperceptible statistical signal into generated text, for example, by partitioning the vocabulary into “green” and “red” lists using a secret key and biasing generation toward the green list. A detector then flags LLM-generated text by testing for a statistical overrepresentation of green tokens.
Prior works (Kirchenbauer et al., 2024b; Liu et al., 2024; Zhao et al., 2024; Lu et al., 2024) suggest that watermarking can withstand common post-edit operations such as insertion, substitution, and deletion, fueling interest in real-world deployment (Bartz and Hu, 2023; Tong, 2024).
@@
At the same time, recent studies (Raffel et al., 2020; Cheng et al., 2025; Wu and Chandrasekaran, 2024; Chen et al., 2024; Jovanović et al., 2024; Diaa et al., 2024) question whether current watermarking methods are sufficiently stress-tested. Existing evasion approaches broadly fall into two categories. Query-based attacks (Wu and Chandrasekaran, 2024; Chen et al., 2024; Jovanović et al., 2024) use repeated queries to infer the green token set and then remove the signal, but they typically assume unrestricted access to the victim model and often partial knowledge of the watermarking scheme (i.e., a gray-box setting). In contrast, query-free attacks (Raffel et al., 2020; Cheng et al., 2025) operate in a realistic black-box setting and usually rely on LLM-based rewriting. However, existing query-free methods achieve only limited evasion success and, crucially, often substantially distort the original semantics.
In this paper, to address these limitations and deepen our understanding of watermarking vulnerabilities, we theoretically analyze rewriting attacks from first principles and show that it suffices for a rewriter (an LLM) to reduce the average conditional probability of sampling green tokens by a margin $\delta>0$ to make the detection probability vanish exponentially in $\delta^{2}$. This characterization directly motivates a practical query-free attack, which we call the Bias Inversion Rewriting Attack (BIRA). BIRA leverages this theoretical insight during rewriting by applying a negative logit bias to a proxy suppression set identified via token self-information, thereby reducing the overrepresentation of green tokens.
Empirically, this simple intervention enables BIRA to evade detection across a range of recent watermarking schemes, substantially outperforming prior query-free attacks while better preserving the original semantics. These results suggest that many current watermarks can be removed with minimal effort, underscoring the need for more robust watermarking methods and more rigorous evaluation beyond simple editing-based stress tests.
@@
  • We develop a theoretical analysis of watermark rewriting attacks, deriving a sufficient condition for successful evasion.
  • Building on this insight, we propose BIRA, a simple and practical query-free attack that weakens the watermark signal by applying a negative logit bias to a proxy suppression set during rewriting.
  • We conduct extensive experiments showing that BIRA achieves state-of-the-art evasion rates against recent watermarking methods while maintaining high semantic fidelity.
@@ 2 Related Work
LLM watermarking. Kirchenbauer et al. (2024a) introduced a widely used green–red list watermarking scheme that partitions the vocabulary into green and red subsets and embeds a detectable statistical signal by applying a positive logit bias to green tokens. Subsequent work has improved robustness by enhancing key generation and detection procedures (Kirchenbauer et al., 2024b; Liu et al., 2023a; Zhao et al., 2024; Liu et al., 2024; Lee et al., 2024; Lu et al., 2024) or by more faithfully preserving the original LLM output distribution (Wu et al., 2023). Other lines of research explore sampling-based watermarking approaches (Aaronson and Kirchner, 2022; Hu et al., 2023; Christ et al., 2024). More recently, sentence-level watermarking methods have been proposed to further enhance robustness by embedding signals at the sentence level rather than the token level (Hou et al., 2024; Dabiriaghdam and Wang, 2025).
Watermark evasion attacks. Watermark evasion attacks can be broadly categorized under the threat model into two types: query-based and query-free. Query-based attacks (Jovanović et al., 2024; Chen et al., 2024; Wu and Chandrasekaran, 2024) identify the green token set by issuing a large number of crafted prefix prompts. While effective, this strategy is computationally expensive and has a limited practicality in rate-limited API settings. Additionally, these methods often require knowledge of the watermarking scheme (Wu and Chandrasekaran, 2024) or access to top-$k$ token probabilities of the watermarked model (Chen et al., 2024), which is typically infeasible for proprietary systems.
By contrast, query-free attacks (Kirchenbauer et al., 2024a; Krishna et al., 2023; Cheng et al., 2025; Diaa et al., 2024) operate directly on generated text in a practical black-box setting, typically via LLM-based rewriting. Prior work either fine-tunes an LLM for paraphrasing (Krishna et al., 2023; Diaa et al., 2024) or masks and regenerates high-entropy tokens (Cheng et al., 2025). However, these methods often achieve limited evasion success and, crucially, can substantially distort meaning; without an explicit fidelity constraint, the task becomes degenerate, since one can simply regenerate fresh unwatermarked text with an unwatermarked LLM. In contrast, our method achieves markedly higher evasion rates while maintaining high semantic fidelity, without additional training and under a black-box setup.
@@ 3 Preliminary
Figure 1: Illustration of BIRA. A watermarked LLM typically increases the likelihood of sampling green tokens by adding a positive bias $\gamma>0$ to their logits at each generation step. In contrast, BIRA applies a negative bias $\beta<0$ to a proxy suppression set, thereby suppressing their sampling probability. This inversion lowers the probability of generating green tokens and weakens the watermark signal, enabling the paraphrased text to evade detection.
@@ 4 Method
In this section, we first analyze rewriting-based evasion and derive sufficient conditions under which watermark detection fails (Section 4.1), and then translate this analysis into a practical query-free attack, Bias Inversion Rewriting Attack (BIRA) (Section 4.2). Figure 1 provides an overview.
4.1 Theoretical Analysis of Rewriting Attack
A watermark evasion attack aims to reduce the statistical overrepresentation of green tokens so that the resulting text no longer triggers the detector. We begin with a simple reduction: for any detector whose test statistic is a nondecreasing function of the empirical green-token rate, detection is equivalent to thresholding the empirical green-token rate $\hat{p}(y;\mathcal{W}_{k})$ (Lemma 4.1). Building on this, we show that if an attack can keep the average conditional probability of sampling a green token below the induced threshold by a margin $\delta>0$, then the detection probability decays exponentially in $\delta^{2}$ (Theorem 4.2).
Lemma 4.1.
@@ 4.1 Theoretical Analysis of Rewriting Attack
$Z(y;\mathcal{W}_{k})$ $=h\!\left(\hat{p}(y;\mathcal{W}_{k})\right),$
$\hat{p}(y;\mathcal{W}_{k})$ $=\frac{1}{N}\sum_{n=0}^{N-1}\mathbf{1}\{y^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\},$
@@
Lemma 4.1 shows that, for this class of detectors, watermark detection depends only on whether the empirical green-token rate $\hat{p}(y;\mathcal{W}_{k})$ exceeds a threshold $p_{\tau}$. We next show that if a rewriter (an LLM) can keep the average conditional probability of sampling a green token at least $\delta>0$ below this threshold across the sequence, then the detection probability decays exponentially in $\delta^{2}$ (Theorem 4.2).
Theorem 4.2.
Let $\tilde{y}=[\tilde{y}^{(0)},\ldots,\tilde{y}^{(N-1)}]$ be the rewriter’s output, and let $p_{\tau}$ denote the induced green-rate threshold from Lemma 4.1. If there exists $\delta>0$ such that
$\frac{1}{N}\sum_{n=0}^{N-1}\mathbb{E}\!\left[\mathbf{1}\{\tilde{y}^{(n)}\in\mathcal{G}(\mathcal{W}_{k})\}\,\middle|\,\tilde{y}^{0:n-1}\right]\;\leq\;p_{\tau}-\delta,$
@@
$\Pr\!\big[\mathcal{D}(\tilde{y},\mathcal{W}_{k})=1\big]\;\leq\;\exp\!\left(-\frac{1}{2}\,N\,\delta^{2}\right).$
Theorem 4.2 shows that if a rewiter can keep the average conditional probability of sampling a green token at least $\delta$ below the detector threshold $p_{\tau}$ across the sequence, then the detection probability decays exponentially in $\delta^{2}$. Thus, even a small per-step suppression of sampling the green-token probability, when achieved on average across the sequence, is sufficient to drive the overall detection probability toward zero. Proofs of Lemma 4.1 and Theorem 4.2 are provided in Appendix B.
@@
Algorithm 1 BIRA (Bias-Inversion Rewriting Attack)
1:Input: System prompt $S$; Watermarked text $\hat{y}^{0:N-1}$; Language model $\mathcal{M}$; Percentile $q\in[0,1)$; Initial bias $\beta_{0}<0$; $\mathrm{lr}>0$; Max restarts $R$; Max length $L_{\text{max}}$; Window size $h$; threshold $\rho\in(0,1]$.
2:Phase 1: Construct Proxy Suppression Set $\widehat{\mathcal{G}}$
3: Compute self-information $I^{(n)}$ for each token $\hat{y}^{(n)}$ using the language model $\mathcal{M}$.
4:for $n=0$ to N1N-1 do
5:  $I^{(n)}\leftarrow-\log P_{\mathcal{M}}(\hat{y}^{(n)}\mid\hat{y}^{0:n-1})$
@@
8: Define the proxy suppression set $\widehat{\mathcal{G}}\leftarrow\{\operatorname{id}(\hat{y}^{(n)})\mid I^{(n)}\geq\eta,\;n\in[0,N-1]\}$
9:Phase 2: Perform Bias-Inversion Rewriting
@@
21:   $\beta\leftarrow\min(0,\beta+\mathrm{lr})$ (Reduce strength of bias and restart)
@@
24:   return $\tilde{y}$ (Return text)
@@ 4.2 Bias-Inversion Rewriting Attack
To translate Theorem 4.2 into a practical attack, one would ideally suppress tokens in the true green set $\mathcal{G}(\mathcal{W}_{k})$, but $\mathcal{G}(\mathcal{W}_{k})$ is inaccessible in a black-box setting. However, Theorem 4.2 implies that perfect identification is unnecessary: evasion only requires a small but consistent reduction in the probability of sampling green tokens. We therefore construct a global proxy suppression set $\widehat{\mathcal{G}}$ from the input text to target likely watermark traces during rewriting.
Although the true green list $\mathcal{G}(\mathcal{W}_{k})$ is unknown, the watermarked text $\hat{y}$ contains the realized watermark trace: specific token choices fixed at generation time, and detection relies on these traces persisting under rewriting. In contrast, unwatermarked generation, produced without access to the secret key, does not systematically bias toward green tokens beyond the baseline rate. Motivated by this, we treat a subset of tokens in $\hat{y}$ as a proxy for watermark traces and suppress their survival during rewriting to reduce green-token bias in the rewritten text.
A naive proxy is to include all input tokens in $\widehat{\mathcal{G}}$, but this can degrade text quality by suppressing frequent function words. Instead, we target tokens most likely to carry watermark influence. Prior work (Cheng et al., 2025) shows that watermarking schemes concentrate their effect on high-entropy positions to preserve utility, making high-surprisal tokens more likely to reflect watermark-biased choices. This is consistent with observations that watermarking is harder to embed in low-entropy settings such as code generation, where limited flexibility to bias token choice leads to fewer green tokens (Lee et al., 2024). We therefore construct $\widehat{\mathcal{G}}$ using token self-information, focusing suppression on likely traces while better preserving text quality.
Given a watermarked text $\hat{y}=[\hat{y}^{(0)},\ldots,\hat{y}^{(N-1)}]$, we compute token surprisal under a public language model $\mathcal{M}$:
$I^{(n)}\;=\;-\log P_{\mathcal{M}}\!\left(\hat{y}^{(n)}\mid\hat{y}^{(0:n-1)}\right).$
Let $\eta$ be the $q$th percentile of $\{I^{(n)}\}_{n=0}^{N-1}$, and define
@@
where $\mathrm{id}(\cdot)$ maps a token to its vocabulary index. During rewriting to produce $\tilde{y}$, we apply bias inversion by adding a negative logit bias $\beta<0$ to tokens in $\widehat{\mathcal{G}}$ at every decoding step:
$l^{(n)}_{u}\;\leftarrow\;l^{(n)}_{u}+\beta\,\mathbf{1}\{u\in\widehat{\mathcal{G}}\},\qquad\forall u\in\mathcal{V}.$
This discourages reproduction of watermark traces, and empirically, suppressing tokens in $\widehat{\mathcal{G}}$ suffices to reduce the average green-token sampling probability (Theorem 4.2).